### v1

# Chasing Anharmonicities in Polarization-Orientation Raman Spectra of Acene Crystals with Machine Learning

Abstract
We present a first-principles machine-learning computational framework to investigate anharmonic effects in polarization-orientation (PO) Raman spectra of molecular crystals, focusing on anthracene and naphthalene. By combining machine learning models for interatomic potentials and polarizability tensors, we enable efficient, large-scale simulations that capture temperature-dependent vibrational dynamics beyond the harmonic approximation. Our approach reproduces key qualitative features observed experimentally. We show, systematically, what are the fingerprints of anharmonic lattice dynamics, thermal expansion, and Raman tensor symmetries on PO-Raman intensities. However, we find that the simulated polarization dependence of Raman intensities shows only subtle deviations from quasi-harmonic predictions, failing to capture the pronounced temperature-dependent changes that have been reported experimentally in anthracene. We propose that part of these inconsistencies stem from the impossibility to deconvolute certain vibrational peaks when only experimental data is available. This work therefore provides a foundation to improve the interpretation of PO-Raman experiments in complex molecular crystals with the aid of theoretical simulations.
I Introduction
Molecular crystals are a class of solids where the molecular units, which constitute the basis of the crystal lattice, are kept together by intermolecular (non-covalent) interactions. The weak nature of these forces, when compared to those responsible in covalent bonds, allows one to identify two distinct energy scales in the nuclear lattice dynamics. At higher energies, intramolecular vibrations usually give rise to phonon branches with flat dispersion, as a result of the weakly perturbed internal motion of the individual molecules. In contrast, at low energies, the collective translations and hindered rotations of the molecular units are responsible for the unique spectral fingerprints of a given crystal structure.
These low-frequency modes, typically below 200 cm-1, dictate numerous properties of organic molecular crystals. It has been discussed, for example, how this region determines polymorph ordering at finite temperatures, by strongly contributing to the vibrational free energy [1, 2, 3], or how it dominates heat [4, 5] and charge [6, 7] carrier mobility in organic semiconductors. Because the weak intermolecular interactions lead to large-amplitude motion, these collective lattice vibrations can exhibit pronounced anharmonic character. Taking into account deviations from the ideal harmonic crystal picture is especially relevant when modeling these systems.
Recently, it was shown [8, 9] that polarization-orientation (PO) vibrational Raman scattering experiments conducted on several molecular crystals, including some linear oligoacenes, reveal an effect that has been attributed to anharmonic coupling between low-frequency intermolecular phonon modes. This effect consists of a deviation of the polarization dependence of the Raman intensity of some phonon modes from the prediction of the harmonic crystal approximation. To rationalize this finding, a two-mode model, allowing phonon-phonon coupling in the description of inelastic light scattering, was presented in Ref. [10]. Parameters that were fitted on this model, based on the experimental data, could successfully explain the pattern that was measured and led to the conclusion that anharmonic mode-coupling terms are necessary for a complete description of the Raman scattering process in molecular crystals.
In this paper, we provide an ab initio quality framework that can be used to reproduce, explain and provide quantitative insights into the type of anharmonic effects that can be directly probed by PO-Raman spectroscopy. As a case study, we focus on linear oligoacene molecular crystals, in particular naphthalene and anthracene. In order to include all orders of anharmonicity of the potential energy surface (PES), our approach is based on the well-established time-correlation function formalism for Raman signals [11, 12]. The standard approach of combining some flavor of ab initio molecular dynamics (AIMD)[13] with density functional perturbation theory (DFPT) [14] to obtain Raman spectra is especially challenging in this case, due to the vast range of experimental conditions one needs to explore, and the large cell sizes required to accurately describe the phonon manifold. Furthermore, to capture subtle variations in the angle-resolved Raman intensity, the spectra have to be thoroughly converged, and finite-size artifacts due to periodic boundary conditions need to be mitigated. Naturally, machine-learning (ML) methods can be exploited in order to make these tasks feasible, but they also need to be scrutinized for accuracy. An array of related ML methods have been previously discussed and benchmarked in the literature [15, 16, 17, 18], but the focus on temperature-dependent PO-signals proves to be a more stringent test of the necessary accuracy of the framework.
In this manuscript, we begin by introducing the theoretical background of Raman scattering, emphasizing the limitations of the harmonic approximation and the need for anharmonic treatments. To address these challenges, we train machine-learning-based models for interatomic potentials and polarizability tensors, benchmarking their accuracy against reference ab initio data. The RGDOS approximation proposed by Komsa and coworkers in Refs. [19, 20] for the polarizability of large supercells is assessed in comparison to more stringent approaches. We then employ these models to simulate the temperature-dependent PO-Raman spectra of anthracene and naphthalene, capturing anharmonic effects related to temperature and the quantum nature of the nuclei. By comparing our results with experimental data, we analyze the polarization dependence of Raman intensities and discuss the subtle deviations from the harmonic predictions. Our findings highlight the interplay between anharmonic lattice dynamics and Raman tensor symmetries, providing insights into the origins of the temperature-dependent behavior of low-frequency Raman modes and suggesting guidelines for the interpretation of the rich amount of information contained in PO-dependent Raman spectra. This work demonstrates the power of combining machine learning with established theoretical frameworks to unravel complex anharmonic phenomena in molecular crystals.
II Theory of Polarization-Orientation Raman Scattering
The Raman effect, first described in 1928 [21], is a process where photons undergo inelastic scattering on matter, resulting in a finite frequency (and polarization) shift between incident and scattered radiation. At a certain energy range, the spectrum of frequency differences is a manifestation of vibrational transitions occurring in the scattering medium, providing unique fingerprints of the vibrations of the material.
Single crystal polarization-orientation Raman experiments are often conducted in a back- (or forward-) scattering geometry, where the incident () and scattered () fields are linearly polarized and propagate through the crystal, which is positioned at a certain orientation [8, 22, 23]. The scattered radiation is then collected with polarization either perpendicular () or parallel () to that of the incident beam. Once the configuration is fixed, the experiment is repeated spanning an angle between and an arbitrary reference, either by rotating the sample or the direction of .
Within a semiclassical treatment, where the light is treated classically and the matter quantum-mechanically, the expression for the Raman cross-section in the electric dipole approximation reads [11, 24]
where is the polarizability tensor of the matter and, within the Born-Oppenheimer (BO) approximation, the summation runs over all vibrational states, with being the thermal equilibrium population associated to the initial state at inverse temperature . Here and are the polarization vectors defined above, either for the parallel or perpendicular setting. The unpolarized Raman spectrum can be obtained by integrating the Raman intensity over all angles in both configurations, and it is defined as
If we restrict our description to non-resonant Raman, i.e. when the energy of the incident radiation is far below any electronic transition of the system, only the static polarizability tensor of the electronic ground state is needed [24]. Within perturbation theory, its components take the form
where is the electronic ground state, the summation runs over all the excited electronic states, is the electric dipole operator and denote Cartesian components. With this definition of the polarizability tensor, Eq. 1 becomes the well known Placzek approximation [24, 25] to the full expression of the light scattering cross section, first derived by Kramers, Heisenberg and Dirac [26, 27]. Note that, as recently highlighted in Ref. [28], by treating the quantity in Eq. 3 as the polarizability tensor (or dielectric susceptibility) entering Eq. 1, we are already excluding the effect of finite scattering wave-vectors, essentially assuming that only vibrations at the -point of the crystal phonon Brillouin zone are probed in the experiment, i.e. , where the are the propagation directions of incoming and scattered fields. This tends to be a very good assumption for first-order Raman scattering, but breaks down in the context of second-order Raman scattering in polar materials [28].
At this stage, the polarizability tensor can be regarded as an operator for the nuclei, and, in the BO picture, it depends parametrically on the nuclear coordinates through the electronic states of Eq. 3. Thus, it is commonly expanded in a Taylor series around the nuclear equilibrium configuration
where runs over all nuclear degrees of freedom. While the first term in the expansion above is responsible for Rayleigh scattering, the higher-order terms, involving polarizability derivatives, enter the expression for the Raman cross-section. If the coordinates are chosen to be the mass-scaled normal modes of the system , these derivatives are referred to as Raman tensors, , i.e., at first order
Note that are Cartesian-space vectors, obtained by removing the mass-scaling that exists in the eigenvectors that come out of the diagonalization of the dynamical matrix. As such, they form a non-orthonormal basis that needs to be handled carefully in common operations such as normal-mode projections of molecular dynamics trajectories, as will be relevant later in the manuscript (see Section S1 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.).
Retaining only the first order term in the expansion of Eq. 4, the matrix element of Eq. 1 is approximated as
At this stage, if the vibrational Hamiltonian is assumed to be harmonic in the normal coordinates, the initial and final states can be written as product states of harmonic eigenstates. It is easy to see, within these two approximations, that the matrix elements of the position operators in Eq. 6 connect only states that differ by a single vibrational excitation in the respective mode, giving rise to the common first-order Raman selection rule.
For the same reason, when the square modulus of Eq. 6 is taken, the cross-products between different eigenmodes vanish, leaving us with the expression for the Raman cross-section
that will depend trivially on temperature through the population of the initial states. These derivations can also be found in the SI of Ref. [10]. We will normally refer to the angular dependence, for a given , as the “PO-pattern” of the corresponding mode in the following.
Eq. 7 is the well known expression for the (harmonic) polarized Raman scattering cross section that is commonly employed to rationalize experimental results in terms of Raman tensors and factor group analysis [30]. As shown above and already clearly stressed in [10], this simple expression is the result of what is often termed a “double harmonic” approximation, as it relies on both harmonic normal modes and the truncation at first order of the expansion in Eq. 4. Clearly, this formula would fall short of accounting for any effect of anharmonic lattice dynamics in the Raman spectrum. Particularly relevant for this work is the fact that in this approximation, the dependence of the (normalized) scattered intensity on the polarization angle is fully determined by the shape, symmetry and relative values of the Raman tensor components of a given mode, and by the geometry of the experiment. Notably, the thermal populations enter only as a multiplicative factor. As the space group of the crystal and the molecular sites do not change with temperature (assuming no phase transitions), the shape and symmetry of are fixed. A common interpretation of this observation is that besides a change in its absolute values, the angular dependence of the intensity calculated with Eq. 7 will be temperature independent in the harmonic picture. As we will see in this paper, though, keeping to Eq. 7 but including temperature-dependent lattice variations can lead to a change in sign and relative magnitude of certain components of , which in turn lead to pronounced variations in the PO pattern that are not connected to dynamical anharmonicity of the vibrational modes.
Any further change with temperature of the PO-patterns must be investigated taking into account dynamical anharmonicity of molecular vibrations, which presents an added challenge to simulations. In particular, dynamical methods which are not based on perturbation theory for the nuclear degrees of freedom, provide an attractive platform to include these effects in their totality. For practical purposes, it is convenient to switch from the energy representation of Eq. 1 to the Heisenberg representation in the time domain, resulting in
where we have conveniently defined
denotes a quantum thermal average, is frequency and is time. Note that, within the BO approximation, the time dependence of the polarizability is given by the time dependence of the nuclear positions. The expression of Eq. 8 can be rewritten as sum over component-wise cross-sections,
highlighting the fourth-rank tensorial nature of , where run over Cartesian components, as recently stressed in both Refs. [10] and [28].
The time-correlation function appearing in Eq. 8 can be well approximated in many situations by considering a classical time-evolution for the nuclei, especially at higher temperatures or when focusing on low-energy vibrational modes. A full quantum-mechanical evaluation of these correlation functions for large and complex systems is too cumbersome with available methodologies, which require the use of explicit wave-function methods or real-time path integrals [31, 32]. However, these expressions are also amenable to approximate methods that join quantum-statistics with classical nuclear dynamics, based on ring-polymer molecular dynamics [33, 34, 35, 36, 37] and centroid molecular dynamics [38, 39, 40, 41]. These methods provide a way to incorporate zero-point-energy and nuclear delocalization effects in the time-correlation functions, but disregard dynamical quantum coherence of the nuclear wavepackets. When coherence effects are small, for example at moderate to high temperatures and in the condensed phase [42], these methods become suitable for including nuclear quantum effects in molecular crystals.
III Simulating PO-Raman Signals
In this work, we focus on capturing, through first-principles simulations, anharmonic effects in the PO-Raman signal coming from low-frequency vibrational modes on polyacene molecular crystals. Going beyond the harmonic approximation without relying on perturbative expansions amounts to evaluating Eq. 8, where the time-dependence of the polarizability tensor through the nuclear coordinates is commonly taken into account by means of AIMD simulations.
Achieving first-principles accuracy in this evaluation relies on being able to obtain calculations of forces and polarizability tensors, through the realization of AIMD trajectories and perturbation theory calculations, resulting in a very high computational cost [43]. Previous studies [1, 44, 45] have used AIMD to successfully compute anharmonic Raman spectra of molecular crystals, albeit computing very few trajectories of the order of 10 ps, resulting on a high uncertainty in peak intensities. In the present study, converged peak intensities are necessary in order to compare polarization-orientation patterns of different temperatures.
A practical and efficient solution is achieved by evaluating the quantities needed in Eq. 8 with highly accurate ML models. In this paper we use two models: one model targets the prediction of energies and forces, needed for the evaluation of normal modes or nuclear time propagation, and the other model targets the prediction of polarizability tensors, needed for the evaluation of the Raman cross sections. When combining ML models in order to calculate a particular physical observable, one needs to take into account that errors from both models will compound in the calculation. This calls for a stringent evaluation of the accuracy and performance of the framework.
Because the training and usage of machine-learning interatomic potentials (MLIPs) is nowadays common in the literature, we only cover the basic aspects of the models used in this work in Methods Section VI.1. Below, we go into more detail on the models used for the polarizability tensors.
III.1 Machine Learning Models for Polarizability Tensors
The polarizability tensor is a second-rank tensor with an isotropic component that transforms as spherical harmonics of and an anisotropic component that transforms as spherical harmonics of . As such, symmetry-adapted equivariant ML methods can deliver a noticeable boost in efficiency of training. All models we discuss in this paper account for the symmetry of the tensor in the learning problem, being thus equivariant. We train two such models here, which are based on an atomic decomposition of this quantity, i.e.
where is the isotropic, traceless part of the polarizability tensor, the anisotropic one and runs over all atoms that define the center of the atomic environments .
The first method we employ to learn this tensor is a MACE model [46, 47], which is an equivariant message-passing graph neural-network (NN). Due to its equivariant formulation, it can be seamlessly extended to predict vectorial and tensorial properties[48, 49], allowing one to set the tensorial rank of the prediction and the symmetries of its components. We will refer to this model as MACE-.
The other method is the symmetry-adapted Gaussian process regression (SA-GPR) proposed in Ref. [50]. As the name proposes, this method is not based on a NN architecture, but instead on a Bayesian Gaussian process regression [51]. This technique casts the problem of learning the target quantity as a problem of finding the coefficients of a linear expansion of the target quantity on a combination of similarity-kernel functions [52]. It is possible to define different kernels corresponding to the different rotational symmetries of each component of , leading to an equivariant “symmetry-adapted” version of the learning problem. We will refer to this model as SA-GPR-.
Both methods mentioned above have already been successfully used in the context of predicting polarizability tensors that were then used to calculate standard vibrational Raman spectra[15, 16, 49, 53, 48, 18]. Here we assess them for the more stringent test of predicting the PO-pattern of each Raman peak.
III.2 The RGDOS Approximation to Polarizability Tensors
The RGDOS approximation to the polarizability tensor, as introduced in Ref. [19] and extended for molecular dynamics in Ref. [20], is an efficient approach to reconstruct the polarizability tensor of a large supercell as a weighted sum of polarizabilities of the multiple unit cells that constitute it. In Ref. [16], Berger et al. showed the good performance of this approximation in reproducing DFT-quality polarizabilities for large supercells and yielding good Raman spectra, including second-order and resonant contributions. The method relies on the existence of a unique equilibrium reference structure and its corresponding eigenmodes, being therefore unable to describe different phases of a material within a single model, but performing well when no phase transitions occur.
Following Ref.[16], one expands the polarizabiliy of the supercell as
where is an atomic displacement vector in the supercell, is the order of the expansion term, is the primitive-cell phonon branch index and is the phonon wavevector, is the corresponding -th order Raman tensor and is the projection coefficient of the supercell displacement onto the primitive-cell eigenmode, as defined in Section S1 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. The most important approximation involved in the expression above is that -th order cross-derivative terms involving different phonon modes are disregarded, while they would be present in an exact expansion of the polarizability tensor for anharmonic systems.
The flexibility of the method stems from the fact that the expansion can be truncated at any desired order. If one is interested only in first-order Raman scattering processes, the task is considerably simplified. In addition to keeping only the first order in the expansion of Eq. 12, only modes at the vicinity of the -point of the Brillouin zone contribute, due to photon momentum conservation. This is well approximated by keeping only contributions from -point phonon modes, i.e.
Once a phonon calculation is performed with the primitive cell and the corresponding Raman tensors are obtained, the -eigenmodes of the supercell are reconstructed by tiling, i.e. concatenating primitive-cell eigenmodes up to the supercell dimensionality. The projection coefficients are then computed at every time-step to yield the polarizability time series for the supercell according to Eq. 13. In the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. we provide more details about how we obtain these projection coefficients in a non-orthogonal basis. Note that a different calculation (phonons + Raman tensors) has to be performed at every temperature if lattice thermal expansion is taken into account, as the equilibrium reference structure will be affected by the change in lattice vectors. The RGDOS PO-Raman spectrum is then obtained by plugging Eq. 13 into Eq. 8, yielding
where anharmonic effects are included in the correlation function of the projected supercell displacements. A further possible simplification regards the truncation of Eq. 13 with respect to the number of phonons considered. The evaluation of Eq. 14 then requires fewer Raman tensors and projection coefficients. We demonstrate in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S2) how retaining only the intermolecular modes in the expansion has no effect in the resulting low-frequency spectrum of the polyacene crystals studied here, making this method particularly attractive for the investigation of low THz regions of molecular crystals.
IV Results and Discussion
IV.1 Performance of MLIPs and Polarizability Models
In the following, we will compare different models for MLIPs and polarizability tensors of polyacene molecular crystals. Berger et al. have performed benchmarking of similar frameworks for different classes of inorganic materials[16]. Here, we apply this approach to the realm of molecular crystals and include a comparison with a state-of-the-art equivariant neural network model for polarizabilities, with the ultimate goal of obtaining ab initio quality PO-Raman spectra.
We start by assessing the quality of the MLIPs we trained for a crucial quantity in the simulation of Raman spectra, namely the frequencies of the crystal vibrations. A particularly challenging region, where we must guarantee that the errors are acceptable, is the low-THz part of the spectrum, which is determined by weak intermolecular interactions often poorly described in MLIPs that are inherently local. Even though we will employ these potentials in dynamical simulations, we choose the harmonic phonons at the -point of the Brillouin zone as a deterministic measure of the quality of our potentials, since they are free of any statistical uncertainties. As detailed in Methods Section VI.1, we compare a second-generation Behler-Parrinello high-dimensional neural network potential [54], which we label HDNNP-MLIP and a equivariant graph neural-network MACE [46] interatomic potential, which we label MACE-MLIP.
We take the anthracene molecular crystal as a benchmark case for evaluating the quality of the machine-learned models for the PES. Details of the training dataset and procedure are given in Section VI. As shown in Fig. 1a MACE-MLIP outperforms the HDNNP-MLIP in the prediction of -point harmonic phonon frequencies resulting in a lower average percentage error of 0.2% against the 1.6% obtained from HDNNP-MLIP. Moreover, MACE-MLIP exhibits a significantly tighter distribution of errors, with a low median and minimal variability, highlighting the consistency and accuracy of this model. In contrast, HDNNP-MLIP shows a broader error distribution, indicating less reliable predictions. When focusing specifically on the low THz region, the error on the harmonic frequency of Raman active modes reported in Fig. 1b is almost negligible for the MACE-MLIP model, while the HDNNP-MLIP largely underestimates the harmonic frequency of the three highest frequency librational modes, with errors up to 20%. These conclusions are consistent with the broader benchmark on accuracy of vibrational properties of ML potentials reported in Ref. [17].
Next, we assess the quality of our ML models for the prediction of the polarizability tensors. In Fig. 2 we show correlation plots for each component of the polarizability tensor of the anthracene crystal, comparing the predicted quantities with the calculated DFPT values over a separate test set of 150 structures (not used for training).
The data shown in Fig. 2 confirms that both methods are very accurate in predicting these tensors, as indeed previous work in the literature has shown [15, 16, 49, 53, 48, 18]. However, we do observe that for the same training dataset size, the MACE- predictions are almost a factor two more accurate than the SA-GPR predictions for diagonal and off-diagonal components of the tensor. One could make the SA-GPR model more accurate by adding more training data, but it is also important to highlight that MACE, as a parametric ML model, does not result in more expensive predictions when increasing the training data set, while SA-GPR, being a non-parametric model, does. This distinction makes MACE the most suitable solution for achieving higher prediction accuracy, especially when applied to larger datasets and system sizes.
In the following, we will use only our two MACE models for the forces and the polarizability predictions for both anthracene and naphthalene.
IV.2 Quality of the Approximations for PO-Raman Signals
Before diving into the analysis of the temperature dependence of the PO-Raman spectra, we analyze the differences between (i) evaluating Eq. 8 with the MACE-MLIP potential for performing the nuclear time evolution combined with the MACE- model for the polarizability evaluations and (ii) evaluating the RGDOS approximation in Eq. 14 with the same MACE-MLIP potential for the nuclear dynamics and fixed Raman tensors computed directly from DFPT. In the following, we name the first procedure “full-ML” and the second “RGDOS-ML”.
In all the calculations performed in this work, the incident and scattered light are polarized in the plane and propagate along the direction in the lab-frame. This means that only the and components of the response are relevant. The PO-Raman spectra we show in the main section are computed in the parallel scattering configuration, as defined in Section II.
In Fig. 3a and b, we show the PO-Raman spectra of the anthracene crystal at 100 K in the range of 0 to 800 cm-1, obtained with full-ML and RGDOS-ML, respectively. All the spectra were computed from supercell trajectories, which amounts to more than 3000 atoms in the case of anthracene. More details are given in Section VI. The overall visual comparison makes it clear that the differences between these two methods are almost imperceptible in this frequency range, with both yielding very similar polarization dependence of all peaks, and following a very similar relative-intensity pattern when regarding different peaks.
The small discrepancies that arise are shown in more detail in Fig. 3c, where the difference between the PO-Raman spectra of anthracene at 100 K obtained with each procedure is shown. Negative values correspond to features that are more intense in full-ML than in RGDOS-ML and vice-versa. In Fig. 3d we report the same difference spectrum, this time for the unpolarized Raman spectrum, as defined in Eq. 2. In general, RGDOS-ML demonstrates remarkable accuracy for PO-Raman analysis in the low-THz region of interest, with discrepancies in the order of of the maximum amplitude. The larger discrepancy is found in the mode of anthracene located below 100 cm-1, where RGDOS-ML overestimates the unpolarized peak amplitude 12%. Consequently, as our focus is limited to first-order Raman – which contains contributions almost solely from the vicinity of -point phonons – we opted to proceed with this framework due to its efficiency and additional advantage of removing uncertainties in the prediction of polarizability tensors.
IV.3 Temperature Dependence of PO-Raman Signals
In Fig. 4 we report the temperature evolution of the parallel configuration PO-Raman spectra of anthracene and naphthalene in the intermolecular motion range of 10 to 170 cm-1, obtained with the RGDOS-ML method. The perpendicular configuration PO-maps can be found in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S3).Both materials crystallize in the same space group, with two molecules per primitive cell, occupying the same sites in both crystals. Factor group analysis thus leads to the same symmetries for their Raman active modes. Namely, the and modes are respectively the three in-phase and the three out-of-phase librational modes of the two molecular units in the primitive cell, and are Raman active[55, 56]. From the three out-of-phase translational modes, two bear and one bears symmetry, being instead IR active. In our simulations, the crystal orientation is along the crystallographic plane (001), therefore, only the components of the Raman cross-section will be measured.
Focusing on panels a and b of Fig. 4, we can see that for both crystals the peaks corresponding to the modes present a four-fold symmetry in their polarization dependence. On the other hand, modes present a larger variation in their PO patterns. Some modes, like and in anthracene still show a four-fold pattern, but with two maxima of different intensities, while other modes, like in anthracene and in naphthalene show a two-fold oscillation pattern. Also, the mode of naphthalene has a weak PO dependence in the parallel configuration, while the peak has zero intensity in the perpendicular experimental geometry (see Section S3 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). These observations are intimately connected to the relative magnitudes of the components of the Raman tensors of the respective modes, which will be discussed in detail in Section IV.3.2.
In panels c–f of Fig. 4, a clear redshift and broadening of all peaks is observed as the temperature increases in both crystals. The extracted peak positions are listed in Table 1. The shift is similar across different modes, spanning a range of cm-1. This effect is a direct result of the incorporation of anharmonicity, in particular the temperature-dependent lattice expansion. As shown in Section S4 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. for anthracene, the redshift is almost negligible when lattice expansion is not taken into account, despite the anharmonic dynamics. This highlights the strong influence of lattice parameter variations on the dynamics of low-frequency intermolecular modes. Notably, some modes even exhibit a blue shift in the absence of lattice expansion. We will analyze the temperature dependence of these PO-patterns in more detail in the following sections, after scrutinizing different sources of PO-pattern modification.
| Anthracene [cm-1] | Naphthalene [cm-1] | ||||||
| Mode | 100K | 220K | 295K | Mode | 100K | 220K | 295K |
| 42 | 38 | 36 | 51 | 45 | 41 | ||
| 52 | 47 | 44 | 58 | 50 | 48 | ||
| 70 | 68 | 67 | 79 | 75 | 69 | ||
| 81 | 77 | 74 | 84 | 79 | 72 | ||
| 121 | 117 | 112 | 115 | 108 | 103 | ||
| 132 | 128 | 124 | 134 | 125 | 119 | ||
IV.3.1 Accounting for nuclear quantum effects
Low-frequency vibrations such as the ones we have discussed in the previous sections are typically well-modeled by classical nuclear motion. The reason is that thermal energy is larger than the energy-spacing between any two vibrational levels of these modes, even at low temperatures ( K cm-1), making classical Boltzmann statistics a good approximation in these cases. In addition, the effective mass of the modes is also large, making classical dynamics equally appropriate at the temperatures (100 to 295 K) and frequency-ranges (20 to 150 cm-1) we are concerned with.
However, mode coupling between these low-frequency intermolecular vibrations and higher frequency modes at the region of cm-1 are known to be present in these crystals [57, 58]. Because these higher-frequency modes are more sensitive to nuclear quantum effects, this coupling can lead to non-trivial temperature-dependence that is sensitive to the quantum nature of the nuclei [59, 60].
In order to investigate these effects, we trained a path-integral coarse-grained (PIGS) MACE MLIP, following Refs.[61, 41]. The elevated temperature was chosen to be 500 K. We performed “elevated-temperature” centroid molecular dynamics (-CMD) [41, 62] simulations for the PO Raman spectrum of anthracene at 100 K and 295 K using the -PIGS method [41]. The results are shown in Fig. 5. As can be seen by comparing the PO-maps in Fig. 5a) and c) with the ones presented in Fig. 4a) and e), the polarization patterns in the intermolecular vibrations region are unaltered by nuclear quantum effects. At higher frequencies, changes are observed. Comparing classical and quantum unpolarized spectra, as shown in Fig. 5b-d, a typical red-shift induced by the inclusion of zero-point motion is observed already at the lower-frequency intramolecular vibrations around 750 cm-1 (highlighted in the insets of panels b-d). Importantly, the temperature dependent changes of the low-frequency modes are not altered by the inclusion of NQE.
While cross-correlations between different vibrational modes could be affected by NQE also in the low-frequency region, these play a minor role in the analysis of PO-Raman spectra, as detailed below. Therefore, we conclude that classical-nuclei simulations are sufficient to obtain anharmonic PO-Raman signals of these molecular crystals at the low-frequency range.
IV.3.2 Changes to PO-Raman Patterns Within a Quasi-Harmonic Approximation
As a first step towards the analysis of the impact of temperature on the polarization dependence of the Raman intensity, we compute the harmonic PO-dependence (parallel configuration) of the Raman intensity in anthracene and naphthalene with Eq. 7, accounting for thermal lattice expansion at 100 K and 295 K, as shown in Fig. 6 for the and modes as an example case. This approach is commonly referred to as quasi-harmonic approximation. For completeness, we show the patterns for all modes in parallel and perpendicular configurations in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S6). In this subsection, when we refer to different temperatures, we are simply referring to the corresponding different lattice constants.
The polarization pattern, i.e the angular dependence of the Raman intensity for a given mode in Eq. 7, is fundamentally determined by the shape, symmetry, and the relative magnitude and sign of the components of the corresponding . We provide explicit expressions for the general angular dependencies of the Raman intensity patterns generated by and tensors in Section S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. In the case of naphthalene and anthracene, with the crystals oriented in the crystallographic plane (001), we show how, once the overall intensity is normalized at each temperature, no variations with lattice expansion can occur in the polarization patterns of modes, both in the parallel and perpendicular configuration within the harmonic approximation of Eq. 7. The same is true for modes in the perpendicular configuration. Only modes in the parallel configuration can be affected by lattice expansion when the relative value of the Raman tensor components of these modes is altered. As a note, we demonstrate in the SI (Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.), that even small inaccuracies in the normal modes can lead to Raman tensors that yield incorrect polarization dependencies. These changes may be overlooked when the variable is integrated out, making PO-patterns a much stricter test for the quality of the theoretical framework compared to simply computing unpolarized spectra.
We start by discussing the PO-patterns of the anthracene crystal. Like in the fully anharmonic case shown in Fig. 4, all modes show a four-fold oscillating pattern across polarization angles, as exemplified in Fig. 6a, and the modes show more variability (see also Fig. S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). Specifically, and have and elements of opposite sign, resulting in a four-fold pattern with two maxima of different intensities, whereas displays a two-fold oscillation due to having these diagonal elements of the same sign. This is an important observation, as despite the fact that a square modulus appears in Eq. 7, the relative sign of the Raman tensor components is qualitatively relevant in the resulting PO-dependence of modes, and can therefore be extracted from a fit of the experimental data. Variations of the PO-patterns at different temperatures are minimal. As a matter of fact, we observe a change in the relative values of the diagonal components of the mode, resulting in a change of intensity of the relative maximum at 270∘ in Fig. 6a.
In the case of naphthalene, the same analysis holds true for the modes, which show a four-fold oscillating pattern, as expected from the fact that these crystals belong to the same space group. The modes again show either a two-fold or a four-fold pattern with maxima of different intensities, depending on the sign of the diagonal components of their Raman tensors (see Fig. S7 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). For example, an inspection of the Raman tensor of mode of naphthalene at 100 K explains both the weak PO-dependence of this mode in the parallel configuration that we observed in Fig. 4 and the absence of PO-dependence in the perpendicular configuration (see Section S3 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.), as the and components are almost equal in magnitude and sign.
Quite surprisingly, though, in naphthalene we observe pronounced changes in the overall polarization patterns of modes with thermal lattice expansion. This is particularly evident in the case of mode , where a significant change in the relative values of the diagonal components of the Raman tensor leads to a transition from a four-fold pattern with two maxima of different intensities at 100 K to a two-fold pattern at 295 K (see Fig. 6b). This observation highlights that even within the quasi-harmonic approximation, temperature-induced lattice expansion can lead to substantial modifications in the polarization dependence of certain Raman active modes. In Section S6 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., we show, by comparing the normal modes displacement vectors at the two different temperatures, that the and modes of naphthalene present a measurable change in direction with the different lattice parameters, leading to significant variations in their Raman tensor components and consequently in their polarization patterns. This is not the case for anthracene, where the normal modes obtained for the structures at different temperatures overlap almost perfectly with each other.
In summary, once the experimental configuration is fixed, only the PO-dependence of modes of a certain symmetry can be affected by thermal lattice expansion in the quasi-harmonic approximation. In the case of anthracene and naphthalene, these are the modes in the parallel configuration. The extent to which the PO-pattern of these modes is affected varies significantly between different materials, as exemplified by the comparison between anthracene and naphthalene. Therefore, while we prove that not all temperature-dependent changes in PO-Raman pattern stem from dynamical coupling between different vibrational modes, we also provide a clear baseline for the analysis of temperature-dependent spectra. The changes that we can capture within the quasi-harmonic approximation, although pronounced, can still be fitted by a second-rank tensor of fixed symmetry. Any additional changes in the polarization patterns must originate from anharmonic dynamical effects involving the vibrational degrees of freedom.
IV.3.3 The Impact of Mode Mixing and Mode Broadening on Anharmonic Spectral Weights
In order to start addressing the role of anharmonicity in the vibrational dynamics of anthracene, we analyze the vibrational density of states (VDOS) and the cross-VDOS (CVDOS) computed in the normal mode basis for the -point intermolecular modes. These quantities are defined in Eq. 16 of Section VI. The VDOS provides a measure of the distribution of vibrational frequencies, while the CVDOS captures correlations between different modes, revealing signatures of mode coupling. One can think of the RGDOS method discussed in Sec. III.2 as a weighting of the -point VDOS and CVDOS with Raman tensors, thus modifying the intensity ratio of the peaks and introducing the angular dependence.
In Fig. 7 we show the VDOS and CVDOS at the point of anthracene at 100 K and 295 K. For completeness, we include the same figure for naphthalene in the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results. (Section S8). At low temperatures, the VDOS of anthracene shows clear Lorentzian-like peaks for each mode. The corresponding CVDOS shows hints of mode coupling between modes -, - and -. The CVDOS of modes with different symmetry is zero and not shown.
As the temperature increases, we observe the emergence of spectral weight in the VDOS of a given mode stemming from other modes. This is particularly evident for the - case at 295 K (inset of Fig. 7b) and provides a clear indication of anharmonic mode mixing. The phenomenon of mode mixing has been addressed in the analysis of Benshalom et al. [10]. They showed that in a two-mode model, the presence of anharmonic coupling leads to significant modifications in the frequency dependence of the Raman response. Instead of isolated Lorentzian peaks, the spectral functions acquire contributions from multiple modes, resulting in nontrivial intensity redistribution and the possibility of temperature-dependent changes in the PO-patterns at a given frequency.
In addition, the broadening of the peaks in VDOS and CVDOS with increasing temperature, evident in Fig. 7, makes it increasingly difficult to assign spectral features to individual modes when one only has access to the PO-Raman maps such as those shown in Fig. 4. This delicate interplay of overlap and mixing between different modes stresses the need for robust deconvolution procedures when analyzing temperature-dependent PO-Raman spectra, as both true anharmonic coupling and spectral overlap can affect the PO-patterns extracted from the fitting procedure of the global maps of Fig. 4.
IV.3.4 Contrasting Simulations with Experimental Data
The first observations of an anomaly in the polarization dependence of the Raman peaks in anthracene and pentacene were made by Asher et al. in 2020 [8]. For example, the polarization dependence of some of the low-frequency Raman peaks of anthracene was found to gradually become less pronounced with increasing temperature. These effects were later reported for other classes of organic molecular crystals by the same authors, and extensive work has been performed in altering (quenching) this behavior through chemical substitutions [9].
This phenomenon, which was described as a dynamical symmetry breaking or mode-specific “liquid-like” transition, has been recently attributed by Benshalom et al. [10] to the manifestation of temperature activated, anharmonic coupling between the intermolecular modes. In their work, the authors tackled the problem from an analytical point of view. An expression for the Raman intensity similar to Eq. 8 was evaluated with fixed Raman tensors in a two-mode model, allowing for anharmonic coupling between modes of same and different symmetries. By relying on fourth-rank tensors for the Raman cross-sections like the ones in Eq. 10, this approach was successful in fitting the experimental PO-dependence of the Raman intensity when the harmonic expression of Eq. 7 failed. These results led the authors to conclude that the observed variation in the PO-dependence of the Raman peaks is a consequence of the anharmonic coupling between intermolecular modes, activated by temperature. Note that the harmonic approximation leading to Eq. 7, as explained in Section II, can only account for deviations due to thermal lattice expansion that conserve the symmetry of the Raman tensors. Wherever this approach fails, this can be regarded as a sign of a dynamical anharmonic effect.
From our point of view, it is instructive to understand whether a first-principles procedure without fitting parameters can capture and provide deeper insight into the observed changes in PO-Raman intensities with increasing temperature. In this study, we are considering the full expression of the anharmonic intensity (i.e Eq. 8), thus retaining the generalization to the fourth-rank tensor of the scattering cross section and the complete mode coupling given by the first-principles full-dimensional anharmonic PES. The expression Eq. 14 has the same form as the starting point of the analysis of Benshalom et al. [10], containing all the cross-correlations terms between phonon modes, i.e. the off-diagonal components of the phonon spectral function, albeit only at the -point. Moreover, our simulations give us the chance to directly compare harmonic and anharmonic results, with and without thermal lattice expansion, providing a tool to decouple and unravel the different effects that temperature has on the resulting Raman spectra.
Before focusing on the angle dependence of the integrated intensity, we highlight some general discrepancies between our simulated and the experimental spectra of anthracene. We observe good overall agreement with the experimental peak positions as reported in Fig. S7 of the SI of Ref. [8], with redshifts in the order of 5-10 cm-1 in the simulated spectra. Notably, the mode has negligible Raman intensity in the experiments, where no polarization dependence could be extracted. In contrast, the mode shows appreciable intensity in our results. This discrepancy is attributed to the level of theory employed, and is not affected by the subsequent dynamics. In fact, it is known to be already present when calculating a harmonic unpolarized Raman spectrum with the same level of theory (PBE+MBD), as reported in Ref. [8].
In this work, we attempt to extract the PO-dependence of the integrated intensity of every peak from the anharmonic PO-maps of Fig. 4 by fitting the Raman spectra of anthracene at 100 K, 220 K and 295 K to a multi-Lorentzian function at each angle. The fitting procedure mimics the one employed for the experiments, as detailed in the SI of Ref. [8], and reported in Section VI. In analyzing the outcome of this procedure, we should keep in mind its shortcomings stemming from the mode broadening effects discussed in Section IV.3.3.
In Fig. 8 we show the fitted PO-dependence of the integrated intensity of the low-frequency Raman peaks of anthracene in both the parallel and perpendicular experimental configurations at 100, 220 and 295 K. The fitted data is normalized to the maximum intensity of each peak at the corresponding temperature. For comparison, we also include the harmonic PO-dependence and the experimental results from Ref. [8], at the closest available temperatures of 80 K, 220 K and 290 K. We remind the reader that for anthracene, the harmonic PO-pattern is insensitive to lattice expansion. In Fig. S10 of the SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results., we show the result of attempting the same fit to the simulated spectra of naphthalene, which was only partially successful for the lowest temperatures, due to the massive broadening at larger temperatures.
The fitted intensity oscillation patterns from the simulated data in the parallel configuration shown in Fig. 8a are in general agreement with the experimental results, with the notable exception of the mode. While both theory and experiment agree that the temperature variation has no impact on the PO-dependence of this mode, its polarization pattern differs. In experiment, it shows a four-fold pattern, with two relative maxima. In the simulations, only a two-fold pattern is present. This discrepancy is entirely determined by the form of the Raman tensor predicted by the specific level of theory employed, and cannot attributed to the subsequent dynamics. Specifically, our PBE Raman tensor has and components of the same sign, which cannot give rise to a four-fold pattern. One must note that experimental artifacts, such as birefringence, can alter the polarization dependence of the Raman intensity in experiment, making the direct comparison even more challenging. Artificially including birefringence in the form of Jones matrices[63] can alter the simulated PO-patterns of modes, but not to a sufficient extent to explain the observed discrepancies (see Section 7 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.). In the perpendicular configuration, all modes present a four-fold pattern in our simulations. This is partially consistent with experiments, where indeed every mode appears as four-fold, but with having two relative maxima of different intensities.
The experimental results in Fig. 8 show a clear change of the PO-dependence with temperature of the and modes in the parallel configuration. The largest change is seen in going from 220 K to 290 K. A similar effect is present for the mode in experiment, but less pronounced. In all cases, the effect can be described as a gradual ”loss” of the oscillation pattern with increasing temperature. When comparing to our simulated, anharmonic PO-maps, we do not find a one-to-one, quantitative correspondence of this phenomenon. In some cases the trend is reproduced correctly, such as the mode, where the oscillation pattern is indeed less pronounced at higher temperatures in both experiment and theory in the parallel configuration. In the case of the and modes, which show a strong loss of polarization dependence in experiment, our simulated pattern is not affected by temperature. Nonetheless, we observe variations of the PO pattern in the mode, with a consistent trend of gradual attenuation of the oscillation pattern with increasing temperature. Given the analysis in Section IV.3.3, we would expect that modes and would both show the effect of temperature due to their coupling, but we only observe changes in . These observations further underline the difficulties of interpreting the fitted PO-Raman maps.
In our simulations, the mode shows variations in the PO pattern with temperature, although we attribute a portion of this variation to the change in the relative magnitudes of the diagonal components of the Raman tensor, discussed in Section IV.3.2 and shown in Fig. 6. In the perpendicular configuration of Fig. 8b, and still show subtle deviations in the simulated patterns. In contrast, the experimental data is substantially less clear regarding the temperature dependence of the PO patterns when the perpendicular geometry is employed. In this experimental geometry, changes are still observed for and , but no monotonic trend can be identified with the increasing temperature.
Our results suggest that, while genuine mode coupling like the one that emerges in Fig. 7 can induce real changes in the polarization dependence, the additional overlap of non-Lorentzian, broadened peaks can also lead to apparent intensity sharing between modes, complicating the interpretation of PO-pattern evolution with temperature. Modes like and , which show a pronounced change in the PO-pattern in experiment, both overlap substantially with an intense peak of a different symmetry: and , respectively. Especially at polarization angles where one of the two overlapping features is in principle not present, any deviation of the other mode spectral feature from a perfect Lorentzian will lead to intensity sharing, thus potentially hiding real anharmonic coupling effects or giving rise to artificial contributions.
V Conclusions
In this work, we have established a computational framework to investigate anharmonic effects in polarization-orientation Raman spectra of molecular crystals, applying it to naphthalene and anthracene. By combining machine learning models for interatomic potentials with the RGDOS approximation, we achieve ab initio accuracy for very large supercells and systematically explore temperature-dependent effects. In particular, we showcased how focusing on the polarization dependence of Raman intensities provides a stringent test of the underlying theoretical framework.
Already at the quasi-harmonic level, we observe that when lattice expansion affects the nature of the normal modes, as in naphthalene, significant changes in the polarization patterns can occur, which are still rationalized within a simplified, second-rank Raman tensor formalism. On top of this, our simulations capture clear anharmonic signatures, including temperature-driven redshifts and peak broadening, which are strongly influenced by lattice thermal expansion. We also confirm that we capture anharmonic mode mixing and that nuclear quantum effects have negligible impact on PO patterns of low-frequency modes. While the polarization dependence of Raman intensities remains largely consistent with harmonic predictions, subtle deviations emerge, reflecting the interplay between anharmonic lattice dynamics and Raman tensor symmetry, complementing the phenomenological approach of Ref. [10].
Notably, even though our approach captures the anharmonic effects on peak positions and intensities, the simulated PO-patterns show only modest temperature dependence compared to the pronounced changes observed experimentally in Ref. [8]. A careful assessment shows that the extraction of PO intensity variations from the full 2D maps that can be gathered from experiments or simulations can be tricky because of the fitting procedure involved. In particular, the simple multi-Lorentzian fitting becomes increasingly ambiguous as temperature-induced peak broadening leads to intensity overlaps between different spectral features. These factors can obscure genuine dynamical effects and complicate a direct comparison between theory and experiment.
Based on the analysis we have presented in this paper, we can propose a way forward. The ability of the RGDOS-ML framework to decompose the PO-Raman patterns into contributions from different vibrational modes in a unambiguous manner could be used to filter intensities in PO-Raman maps and avoid the need to fit the entire pattern at once. More specifically, this could be done by considering the mode-resolved version of Eq. 14 and using it to assign PO-Raman intensities based on the self and cross-correlation terms that result from this analysis. This first-principles decomposition could then be used as a proxy for the analysis of the experimental PO-Raman maps through a more robust global fitting procedure that takes mode mixing and non-trivial spectral overlap into account [10].
It is important to emphasize that a perfect match between simulated and experimental PO-Raman patterns is not expected, as discrepancies already arise at the harmonic level due to the specific level of theory employed for the electronic structure and Raman tensor calculations. Other limitations include the use of finite supercell sizes which impose a finite sampling of the Brillouin zone (here a -grid), and exclude -points in the close vicinity of the Brillouin zone center. While the large supercell still mitigates finite-size effects and indirectly affects the dynamics at the -point, our analysis of the Raman signals is fundamentally restricted to the -point itself, as we stressed from the beginning. We note that allowing for contributions from the vicinity of the -point in Raman spectroscopy could strongly affect the mode-coupling picture, by breaking fundamental symmetry restrictions and allowing modes of different irreducible representations to couple [10]. These contributions are very challenging to assess by molecular dynamics alone, requiring very large system sizes.
Nevertheless, our framework demonstrates that, starting from a given theoretical baseline, we can systematically include temperature and dynamical (anharmonic) effects on top of it, providing a predictive route to assess their impact on the PO-Raman response.
Overall, our results demonstrate that ML-driven simulations combined with established theoretical frameworks can unravel complex anharmonic phenomena in molecular crystals. The insights gained here advance the interpretation of PO-Raman spectra and provide a foundation for future improvements, including more robust deconvolution methods for analysis of 2-dimensional polarization orientation Raman patterns and larger-scale simulations to further bridge theory and experiment.
VI Methods and computational Details
VI.1 Training Machine Learning Models for Interatomic Interactions
In this paper we investigate two types of MLIPs based on neural-network (NN) architectures. One architecture is that of high-dimensional neural-network potentials (HDNNP) [54, 64], which uses two-body and three-body local atomic descriptors in the form of atom-centered symmetry functions. In this representation, there only a few hidden layers and a few nodes on the NN attached to each atomic species are needed, and the construction of the potential is successful for quite complex systems [65, 66, 67, 68, 69]. This architecture does not explicitly encode the equivariance of the transformation of each component of the symmetry functions. The target of the optimization of the NN parameters is the total energy of the system, constructed from the sum of the individual contributions of the atomic energies predicted by the NN, and forces are readily obtained by back-propagation (i.e., the model is conservative). We have referred to this model as HDNNP-MLIP in the main text.
The other architecture is the equivariant message-passing graph NN employed in the MACE model [46, 47]. In this architecture, atoms are represented as nodes of a graph and the interactions between pairs of atoms are the edges. The messages are constructed as many body expansions where each term is approximated by a linear combination of a complete basis of equivariant features [70]. These messages are thus of high-body order and can be understood as an efficient evaluation of a many-body expansion on ACE descriptors [71]. Forces are obtained by automatic differentiation and the model is also conservative. The use of effective high body-order messages delivers highly accurate predictions. The imposition of equivariance reduces the complexity of the problem by incorporating the rotational properties of the representation, resulting in the need of less training data points and removing the need of explicit data augmentation. We have referred to this model as MACE-MLIP in the main text.
In order to construct the training dataset for anthracene, we used the experimentally available lattice parameters at temperatures of 100 K and 295 K from Ref. [8] and performed a 2 ns NVT-MD run with the foundational model MACE-OFF [72] to generate trial structures. A total of 1200 configurations were selected through farthest point sampling and the corresponding total energies and forces were recomputed at the PBE [73] level of theory with many-body dispersion correction (MBD) [74] using the FHI-aims [75] electronic structure package, with the recommended tight basis-set. Dispersion corrections are essential to capture the cohesive forces in molecular crystals and the MBD scheme has proven to be very accurate for structural properties of these systems [76]. Both of our models for anthracene were subsequently trained on this dataset, with 10% of the structures left as test set. The accuracy of the models on training and test set is reported in Table 2.
| Energy (meV/atom) | Force (meV/Å) | |||
| Training | Test | Training | Test | |
| HDNNP-MLIP | 0.3 | 2.5 | 45.6 | 47.7 |
| MACE-MLIP | 0.2 | 0.2 | 5.3 | 7.1 |
| Energy (meV/atom) | Force (meV/Å) | |||
| Training | Test | Training | Test | |
| MACE-MLIP | 0.1 | 0.1 | 2.6 | 3.4 |
For naphthalene, we used the MACE potential decribed in Ref.[17]. The underlying electronic-structure is also PBE+MBD with tight basis-set. For further details on the training procedure, the reader is referred to Section II of Ref.[17]. The errors on energies and forces are reported here (Table 3) for completeness.
VI.2 Training Procedure of the Machine Learning Polarizability Models
The dataset for both polarizability models consisted in 1000 structures sampled from NVT MD simulations generated by our MACE-MLIP. The structures were sampled at 100 K, 150 K and 293 K for anthracene and 30 K, 80 K and 295 K for naphthalene, employing available experimental lattice constants from Refs. [8, 77]. The polarizability tensors were computed from DFPT as implemented in the FHI-aims code using the LDA functional, which has been shown in Ref. [78]to be accurate in capturing the polarizability variations that enter the Raman intensity calculation as long as the underlying dynamics is generated by a high quality potential. We also performed an explicit comparison of LDA vs PBE for Raman tensors in Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results..
VI.3 Phonons calculations and geometry optimizations
For the comparison of Section IV.1 we computed the harmonic phonon frequencies for anthracene on the minimum energy atomic configuration with lattice vectors fixed to the experimental structure at 295K from Ref. [8] with both our ML models and DFT. The geometry optimizations were separately performed with FHI-aims (PBE+MBD, tight basis set) for the DFT case and with our two ML potentials. Phonon calculations were subsequently performed using the phonopy code [81, 82].
VI.4 Raman tensors
The Raman tensors needed for the RGDOS expansion were obtained from finite differentiation of DFPT polarizabilities of displaced minimum energy structures along the normal coordinates. The procedure is repeated for every material at every different lattice parameters considered. The PBE functional was employed in the DFPT calculations with tight basis-sets. Again, a comparison of LDA vs PBE Raman tensors is reported in Section S5 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results..
VI.5 Production simulations for Raman spectra
For the production runs, we performed 48 independent NVE-MD simulations of 100 ps each with a time step of 1 fs for both anthracene and naphthalene at temperatures of 100 K, 220 K and 295 K. For both crystals, we employed supercells to mitigate finite-size effects. This amounts to a total of 3072 atoms for anthracene and 2304 atoms for the naphthalene crystal. The convergence of the PO-Raman spectra with the supercell dimension is reported in Section S9 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. The NVE trajectories were generated from a 2 ns parent NVT trajectory where temperature was controlled with a stochastic velocity-rescaling thermostat [84] with a time constant of 100 fs. The polarizability tensors were computed on-the-fly every 10 fs from the RGDOS expansion Eq. 13, with the procedure to obtain projections coefficients described in Section S1 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. The Raman spectra were subsequently computed through numerical Fourier transformation with a Hanning windowing function applied to the time series. No additional broadening was introduced by the windowing process.
VI.6 Fitting of PO-dependence
Following Ref. [8] we fitted the PO-dependence of the integrated intensity of every peak in the PO-maps, as presented in Figs. 4 and 5 for example, to a multi-Lorentzian function of the form
where is the amplitude, the position and the width of the -th peak. The position and the width for each peak were determined from a first round of fitting at fixed polarization angles at which peaks were as clearly separated as possible. Once determined, under the well-founded assumption that the polarization angle does not influence the peak position and width but only the amplitude, we fitted the oscillating intensities.
VI.7 VDOS and CVDOS calculations
The VDOS and CVDOS of Fig. 7 were computed from Fourier transformation of the velocity autocorrelation function of the -point projections with being a -point normal mode, following the same procedure as Ref. [85]. The final expression reads as
where gives the VDOS of mode and gives the CVDOS between modes and . When defined like this, the VDOS is normalized to twice the vibrational kinetic energy of the system, . The projection velocities are obtained through finite differentiation of normal mode projection coefficients along the MD trajectory. These are defined as
where are the atomic displacements at time and is the cartesian normal mode . In order to extract only the -point projection coefficients, we follow the procedure we described in Section S1 of SM 111See Supplemental Material at [URL will be inserted by publisher] for details on normal modes projections, supercell convergence, perpendicular configuration PO-maps, Raman tensors benchmarking and full harmonic PO-Raman results.. No windowing was applied to the time series before Fourier transformation.
Author contributions
P.L. and M.R. designed and conceptualized the research. S.S. and P.L. trained machine-learning models. P.L. ran simulations, conducted theoretical derivations, analyzed data and produced all figures in the manuscript. M.R. contributed to the data analysis the discussion. P.L. and M.R. wrote the first draft of the paper. All authors edited the draft.
Conflicts of interest
The authors declare no conflicts of interest.
Data availability
Data pertaining to this paper will be made available in a public repository upon publication.
Acknowledgements
M.R. acknowledges funding by the European Union (ERC, QUADYMM, 101169761). P.L. and S.S. acknowledge support from the UFAST International Max Planck Research School. The authors are grateful to Nimrod Benshalom, Noam Pinsk, and Omer Yaffe for insightful discussions, valuable input, and for inspiring this work through their earlier experiments. We further acknowledge Nimrod Benshalom for his critical reading of the manuscript. We also thank Burak Gurlek, Elia Stocco, and George Trenins for helpful discussions.